# Measuring Gravitational Wave Spectrum from Electroweak Phase Transition and Higgs Self-Couplings

### v2

@@ Abstract
In this work, we demonstrate the complete process of using space-based gravitational wave detectors to measure properties of the stochastic gravitational wave background resulting from a first order electroweak phase transition, to infer the parameters governing the phase transition dynamics as well as that of the underlying particle physics model, and eventually to make predictions for important physical observables such as the Higgs cubic and quartic self-couplings which are difficult to measure at colliders. This pipeline is based on a frequency domain simulation of the space-based gravitational wave detector Taiji, taking into account dominant instrumental noises and astrophysical background, where the data analysis is carried out using both the Fisher information matrix and Bayesian inference with Markov-Chain Monte Carlo numerical sampling. We have applied this framework to the simplest extension of the Standard Model, the singlet extension, and show the measured uncertainties of the parameters at various levels of inference, and show that the Higgs cubic and also the quartic coupling can be highly constrained from gravitational wave measurement. We also show the impact from the problem of parameter degeneracy, highlighting the corresponding limitation on parameter inference and on making predictions.
@@ 1 Introduction
The detection and precise measurement of the stochastic gravitational wave background (SGWB) of cosmological origin provide a unique probe of the early Universe (see, e.g., Caldwell:2022qsj; LiGong:2024qmt for recent reviews). Recent results from pulsar timing arrays (PTAs) NANOGrav:2023gor; NANOGrav:2023hde; EPTA:2023sfo; Athron:2023mer have revealed the first tantalizing evidence of a nanohertz SGWB. Ground-based interferometers, including Advanced LIGO LIGOScientific:2017vtl, Virgo Kumar:2024bfe, and KAGRA KAGRA:2020tym, explore the higher-frequency band from tens to thousands of hertz, and their latest observations have placed strong upper limits in the absence of a detection LIGOScientific:2025bgj. Looking ahead, space-based interferometers such as LISA LISA:2017pwj; Robson:2018ifk; LISACosmologyWorkingGroup:2022jok, Taiji Hu:2017mde; Ruan:2018tsw; Wu:2018clg, and TianQin TianQin:2015yph; TianQin:2020hid; Luo:2020bls will extend the accessible frequency window into the millihertz range. Among the most intriguing cosmological sources of the SGWB are those arising from first-order phase transitions (FOPTs) in the early Universe (see, e.g., Caprini:2015zlo; Weir:2017wfa; Mazumdar:2018dfl; Caprini:2019egz; Bian:2021ini; Athron:2023xlk for reviews), which offer a valuable link between cosmology and particle physics by probing symmetry breaking at different energy scales. Of particular interest is the electroweak phase transition (EWPT) associated with the breaking of the electroweak symmetry. In many extensions of the Standard Model, the EWPT can become strongly first order, generating the out-of-equilibrium conditions required for baryogenesis Morrissey:2012db and simultaneously producing a potentially observable GW signal within the reach of future space-based detectors Caprini:2015zlo.
@@
Previous phenomenological studies of the EWPT have mostly relied on simple signal-to-noise ratio (SNR) calculations, combining the predicted SGWB with detector sensitivity curves to assess the detectability of the signal. In practice, however, experimentally extracting a faint cosmological SGWB from the observational data of a space-based detector is particularly challenging due to the presence of many noise sources and the contamination from astrophysical foregrounds (see, e.g., Boileau:2020rpg; Boileau:2021sni).
@@
Space-based GW detectors must contend with various sources of instrumental noise, dominated by laser frequency fluctuations and supplemented by residual acceleration noise, optical-path disturbances, etc LISA:2017pwj. In ground-based experiments, such challenges can be partly mitigated through the use of a detector network, where cross-correlation between independent data streams suppresses uncorrelated noise Allen:1997ad; Romano:2016dpx. Although a network of space-based detectors has been proposed Ruan:2020smc; Cai:2023ywp; Chen:2024fto, most current studies—including the present work—focus on the single-detector case, which requires a different approach for signal extraction. In this context, time-delay interferometry (TDI) Tinto:2014lxa is employed to cancel the dominant laser frequency noise, resulting in three independent data combinations, e.g., the $A$, $E$, and $T$ channels, each with distinct sensitivity to GWs (see, e.g., Smith:2019wny for a recent overview). Among these, the so-called null channel $T$ Adams:2010vc is particularly valuable: at low frequencies, it is almost insensitive to GWs, allowing it to serve as an internal noise monitor that assists in instrument calibration and in distinguishing signal from noise.
@@
In addition, the stochastic signal of cosmological origin is blended with strong astrophysical foregrounds, most notably the unresolved galactic binaries that form a confusion background and the contribution from extragalactic compact binary coalescences Tinto:2004wu; Smith:2019wny; Barack:2004wc; Caprini:2024hue; Boileau:2020rpg. Disentangling these different components—and thereby establishing a confident detection of the SGWB from a FOPT—requires a dedicated statistical analysis.  [^1] A further challenge arises from deterministic, individually resolvable signals such as those from galactic binaries, massive black hole binaries, and extreme mass-ratio inspirals, which must be jointly analyzed along with stochastic components in the so-called global fit framework Cornish:2005qw; Littenberg:2023xpl; Rosati:2024lcs. This formidable task remains under active development and will not be considered in this work.
@@
Significant progress has recently been achieved both in theoretical modeling and in data analysis techniques. The goal of this work is to bridge these two fronts—to integrate state-of-the-art theoretical developments with realistic experimental considerations—and to assess the achievable scientific outcomes, in line with recent efforts within the community Gowling:2021gcy; Gowling:2022pzb; Boileau:2022ter; Caprini:2024hue; Lewicki:2024xan; Huang:2025uer.
@@
For phenomenological studies, the interplay between GW observations and collider measurements has become a major focus of research Ramsey-Musolf:2019lsf. In particular, the cubic and quartic self-couplings of the Higgs boson—central to understanding the scalar potential and the mechanism of electroweak symmetry breaking, but notoriously difficult to measure at colliders DiVita:2017vrr; Zabinski:2023jhr—serve as key observables that highlight the complementarity between collider and GW approaches Alves:2018jsw.
@@
In this work, we demonstrate the above pipeline using the simplest extension of the Standard Model that includes a real scalar singlet—the xSM Barger:2007im; Profumo:2007wc; Profumo:2014opa; Huang:2017jws; Goncalves:2024vkj. This study extends our previous phenomenological analyses Alves:2018jsw by introducing a more realistic statistical treatment tailored to space-based GW experiments.
The remainder of this paper is organized as follows. In Sec. 2, we provide a brief introduction to the xSM and its phase transition dynamics. Sec. 3 introduces the GW spectrum from the EWPT and the associated astrophysical foreground and background models. Sec. 4 describes the detector response, noise modeling, and data simulation procedures. In Sec. 5, we present the statistical inference framework, comparing Fisher forecasts with MCMC results. Sec. 6 applies these methods to the simulated data, mapping spectral constraints onto the xSM parameters and then making predictions for the Higgs self-couplings. Finally, Sec. 7 concludes with a discussion of the implications for future GW missions and collider programs.
@@ 2 The Model
In this section, we introduce the particle physics model and its parameter space, defined by five independent parameters. Once the model parameters are fixed, the phase transition parameters can be determined; these are the quantities that directly enter the prediction of the GW spectrum. Throughout this work, we follow the same convention as in our previous study Alves:2018jsw, to which we refer the reader for further details.
@@
The model “xSM” is the simplest extension of the Standard Model and has been extensively studied in the literature due to its simplicity and its ability to accommodate a strongly first-order phase transition Barger:2007im; Profumo:2007wc; Profumo:2014opa; Huang:2017jws. This model is obtained by coupling the SM Higgs doublet with a real scalar gauge singlet, $S=v_{s}+s$, with the following potential
@@
with its vacuum expectation value $v_{\mathrm{EW}}$ and $G^{+},G^{0}$, the Goldstone modes. Through the two minimization conditions around the vacuum ($v_{\mathrm{EW}}$$v_{s}$), $\mu$ and $b_{2}$ are given by:
@@
These five parameters define the parameter space of the xSM model, which is constrained firstly by the theoretical requirement of the Higgs potential and by experimental measurements of various physical observables. The former includes the boundedness of the potential from below, electroweak vacuum stability at zero temperature, perturbativity, perturbative unitarity, etc. The latter includes precision measurements of the $W$ mass, the oblique electroweak corrections, the Higgs signal strengths, searches for a heavier Higgs-like scalar, etc. These have been extensively studied in the literature and our previous study Alves:2018jsw. Our study in this work is based on the phenomenologically viable parameter space, characterized by a large dataset of parameter space points that pass above the constraints. The readers are referred to the previous work for further details on the phenomenological constraints.
@@
$\begin{split}&i\lambda_{h_{1}h_{1}h_{1}}=6\big[\lambda vc_{\theta}^{3}+\tfrac{1}{4}c_{\theta}^{2}s_{\theta}(2a_{2}v_{s}+a_{1})+\tfrac{1}{2}a_{2}vc_{\theta}s_{\theta}^{2}+\tfrac{1}{3}s_{\theta}^{3}(3b_{4}v_{s}+b_{3})\big],\\ &i\lambda_{h_{1}h_{1}h_{1}h_{1}}=6(\lambda c_{\theta}^{4}+a_{2}s_{\theta}^{2}c_{\theta}^{2}+b_{4}s_{\theta}^{4}).\end{split}$ (2.7)
@@ 2.2 Phase Transition Parameters
For the GW spectrum generated by phase transitions, it is not the set of five model parameters in Eq. (2.6) that appears directly, but rather a derived set of phase transition parameters Caprini:2024hue:
@@
which describes the thermodynamic properties of the EWPT. Specifically, $T_{\mathrm{n}}$ is the temperature at which the phase transition occurs, $\alpha$ quantifies the released vacuum energy normalized to the radiation energy density at $T_{\mathrm{n}}$, $\beta/H_{\mathrm{n}}$ characterizes the inverse time scale of the transition, and $v_{w}$ denotes the bubble wall velocity. These parameters are computed utilizing the most recent version of PhaseTracer with action fitting enabled. Bian:2025yfj; Athron:2024xrh; Athron:2020sbe.
In the Bayesian statistical analysis presented later, we do not directly infer the above thermodynamic parameters. Instead, our analysis is performed on a more direct spectral model characterized by two parameters: the overall amplitude $\Omega_{0}$ and the peak frequency $f_{\mathrm{p}}$. This choice avoids the issue of parameter degeneracy that would arise if one were to analyze the thermodynamic parameters directly, as will be discussed in subsequent sections. Once constraints on $\Omega_{0}$ and $f_{\mathrm{p}}$ are obtained, we then employ the mapping procedure introduced later to translate these results into constraints on the thermodynamic parameters and further into the particle-physics model parameters defined in Eq. (2.6). The detailed relations between these parameter sets, which require the calculation of the finite-temperature effective potential and the associated phase transition dynamics, can be found in our previous work Alves:2018jsw and will not be repeated here.
Finally, we note that there remain significant theoretical uncertainties (see, e.g., Croon:2020cgk; Athron:2022jyi; Athron:2023rfq; Lewicki:2024xan; Zhu:2025pht) in determining the above phase transition parameters. These arise from several sources: the construction of the finite-temperature effective potential and the nucleation rate calculation, including the issue of gauge dependence for the perturbative calculation of the effective potential Patel:2011th; Arunasalam:2021zrs; Lofgren:2021ogg; Hirvonen:2021zej; and the calculation of the parameters $T_{\mathrm{n}}$, $\alpha$, $\beta/H_{\mathrm{n}}$, and $v_{w}$, where different approaches with varying levels of rigor exist Guo:2021qcq. For simplicity, we do not consider these theoretical uncertainties in the following statistical analysis, and we refer to Lewicki:2024xan for a study of this kind based on the FIM analysis.
@@ 3 Gravitational Waves: Cosmological and Astrophysical
While the target is the GW signal from the EWPT, one must also account for astrophysical and other possible cosmological backgrounds or foregrounds. These additional signals contribute incoherently to the observed data, thereby increasing the challenge of detecting the target signal. As a first step and for simplicity, we include the background from extragalactic compact binary coalescences, while a more complete treatment of other additions will follow the same methodology, involving only additional computational complexity.
@@ 3.1 Electroweak Phase Transition
In a first-order phase transition, vacuum bubbles are nucleated, then grow, collide, and merge, driving the universe from a metastable to a stable state Guth:1981uk; Hindmarsh:2020hop. The expansion and collision of these bubbles disturb the surrounding plasma in the form of sound waves and magneto-hydrodynamic (MHD) turbulence Caprini:2015zlo; Hindmarsh:2013xza. As a result, the SGWB from a first-order phase transition has three main contributions: (i) bubble collisions Kosowsky:1991ua; Kosowsky:1992rz; Kosowsky:1992vn; Huber:2008hg; Jinno:2016vai; Jinno:2017fby, (ii) sound waves Hindmarsh:2013xza; Hindmarsh:2015qta; Cai:2023guc; RoperPol:2023dzg, and (iii) MHD turbulence Kosowsky:2001xp; Caprini:2009yp; Binetruy:2012ze; Gogoberidze:2007an; Niksa:2018ofa; RoperPol:2019wvy. However, the bubble-collision contribution is usually negligible compared to sound waves for the EWPT, and that of the MHD turbulence is both subdominant and still uncertain. Thus, we consider only the dominant contribution from sound waves.
The spectrum from sound waves takes the following form Caprini:2015zlo; Guo:2020grp:
@@
$\left(\frac{7}{4+3(f/f_{\mathrm{p}})^{2}}\right)^{7/2}\Upsilon\!\left(\tau_{\mathrm{sw}}\right)\;,$
where $H_{\mathrm{n}}$ is the Hubble rate at the nucleation temperature $T_{\mathrm{n}}$, $f_{\mathrm{p}}$ is the peak frequency Romero:2021kby,
@@
and $\Upsilon$ is a suppression factor Guo:2020grp, a function of the finite lifetime Ellis:2018mja of the sound waves $\tau_{\text{sw}}$, which in a radiation-dominated universe takes the following form:
@@
while the formula for a universe with a generic expansion rate was derived recently in Guo:2024kfk.
The lifetime $\tau_{\text{sw}}$ is usually chosen to be the time for the onset of MHD turbulence, $\tau_{\mathrm{sw}}=\frac{R_{\mathrm{pt}}}{\overline{U}_{f}}\;$ Weir:2017wfa, where $R_{\mathrm{pt}}=(8\pi)^{1/3}v_{\mathrm{w}}/\beta$ is the mean bubble separation, and $\overline{U}_{f}=\sqrt{\frac{3\kappa_{\mathrm{sw}}\alpha}{4(1+\alpha)}}$ is the root-mean-square fluid velocity Weir:2017wfa. For convenience, one can also define $K\equiv\frac{\alpha}{1+\alpha}$ Caprini:2024hue, which encapsulates the dependence on transition strength.
@@
We set the spectral parameters to be log uniformly distributed in the ranges $\Omega_{0}\in(10^{-20},10^{-5})$ and $f_{\mathrm{p}}\in(10^{-5},1)\penalty 10000\ \mathrm{Hz}$, as the priors, taking into account the theoretical constraints of the xSM (as illustrated in Fig. 9) and following also Caprini:2024hue; Romero:2021kby. The corresponding ranges for phase transition parameters are: $\beta/H_{\mathrm{n}}\in(10^{-1},10^{5})$, $T_{\mathrm{n}}\in(1,10^{5})\penalty 10000\ \text{GeV}$, and $K\in(10^{-3},1)$, consistent with the sampling reconstruction results shown in Fig. 8. In addition, the wall velocity and the relativistic degrees of freedom are fixed to $v_{w}=1.0$ and $g_{*}=100$, respectively for simplicity, and the efficiency factor $\kappa_{\text{sw}}$ is taken as a function of $\alpha$ and $v_{w}$ (see the appendix of Espinosa:2010hh for the fitting formula). We note that choosing different values of the parameters such as $v_{w}$ leads to corresponding changes in the Bayesian inference, but not qualitatively on the conclusions obtained in this work.
@@
$v_{s}\simeq 26.62\ \mathrm{GeV},\quad m_{h_{2}}\simeq 852.74\ \mathrm{GeV},\quad\theta\simeq 0.113,\quad b_{3}\simeq 1920.8\ \mathrm{GeV},\quad b_{4}\simeq 4.15,$
@@
$\frac{\beta}{H_{n}}=676.1,\quad T_{\mathrm{n}}=43.65\penalty 10000\ \mathrm{GeV},\quad K=0.394,$
@@
$\Omega_{0}=3.16\times 10^{-12},\quad f_{\mathrm{p}}=5.62\times 10^{-3}\penalty 10000\ \mathrm{Hz}.$

### v3

@@ Abstract
In this work, we demonstrate the complete process of using space-based gravitational wave detectors to measure properties of the stochastic gravitational wave background arising from a first-order electroweak phase transition. Based on frequency-domain simulations of the Taiji mission, including instrumental noise and astrophysical foregrounds, we perform parameter inference using both the Fisher information matrix and Bayesian Markov-Chain Monte Carlo sampling. We show how the reconstructed spectrum constrains the macroscopic parameters of the phase transition, and further how these constraints map onto the underlying particle-physics parameters in a singlet-extended Standard Model. Our results demonstrate that the Higgs cubic and quartic self-couplings can be significantly constrained using gravitational wave observations, despite limitations arising from parameter degeneracy.
@@ 1 Introduction
The detection and precise measurement of the stochastic gravitational wave background (SGWB) of cosmological origin provide a unique probe of the early Universe (see, e.g., Caldwell and others (2022); LiGong et al. (2025) for recent reviews). Recent results from pulsar timing arrays (PTAs) Agazie and others (2023a, b); Antoniadis and others (2023); Athron et al. (2024b) have revealed the first tantalizing evidence of a nanohertz SGWB. Ground-based interferometers, including Advanced LIGO Abbott and others (2018), Virgo Kumar and Dent (2024), and KAGRA Akutsu and others (2021), explore the higher-frequency band from tens to thousands of hertz, and their latest observations have placed strong upper limits in the absence of a detection Abac and others (2025b, a). Looking ahead, space-based interferometers such as LISA Amaro-Seoane and others (2017); Robson et al. (2019); Auclair and others (2023), Taiji Hu and Wu (2017); Ruan et al. (2020a); Wu (2018); Zhang et al. (2024); Deng et al. (2025); Ren et al. (2023); Du and others (2025); Du et al. (2024); Jiang and Huang (2026); Wu and Tang (2025), and TianQin Luo and others (2016); Mei and others (2021); Luo and others (2020) will extend the accessible frequency window into the millihertz range. Among the most intriguing cosmological sources of the SGWB are those arising from first-order phase transitions (FOPTs) in the early Universe (see, e.g., Caprini and others (2016); Weir (2018); Mazumdar and White (2019); Caprini and others (2020); Bian and others (2021); Athron et al. (2024a) for reviews), which offer a valuable link between cosmology and particle physics by probing symmetry breaking at different energy scales. Of particular interest is the electroweak phase transition (EWPT) associated with the breaking of the electroweak symmetry. In many extensions of the Standard Model, the EWPT can become strongly first order, generating the out-of-equilibrium conditions required for baryogenesis Morrissey and Ramsey-Musolf (2012) and simultaneously producing a potentially observable gravitational wave (GW) signal within the reach of future space-based detectors Caprini and others (2016).
@@
Previous phenomenological studies of the EWPT have mostly relied on simple signal-to-noise ratio (SNR) calculations, combining the predicted SGWB with detector sensitivity curves to assess the detectability of the signal. In practice, however, experimentally extracting a faint cosmological SGWB from the observational data of a space-based detector is particularly challenging due to the presence of many noise sources and contamination from astrophysical foregrounds (see, e.g., Boileau et al. (2021a, b)).
@@
Space-based GW detectors must contend with various sources of instrumental noise, dominated by laser frequency fluctuations and supplemented by residual acceleration noise, optical-path disturbances, etc Amaro-Seoane and others (2017). In ground-based experiments, such challenges can be partly mitigated through the use of a detector network, where cross-correlation between independent data streams suppresses uncorrelated noise Allen and Romano (1999); Romano and Cornish (2017). Although a network of space-based detectors has been proposed Ruan et al. (2020b); Cai et al. (2024); Chen et al. (2025), most current studies—including the present work—focus on the single-detector case, which requires a different approach for signal extraction. In this context, time-delay interferometry (TDI) Tinto and Dhurandhar (2014) is employed to cancel the dominant laser frequency noise, resulting in three independent data combinations, e.g., the $A$, $E$, and $T$ channels, each with distinct sensitivity to GWs (see, e.g., Smith et al. (2019) for a recent overview). Among these, the so-called null channel $T$ Adams and Cornish (2010) is particularly valuable: at low frequencies, it is almost insensitive to GWs, allowing it to serve as an internal noise monitor that assists in instrument calibration and in distinguishing signal from noise.
@@
In addition, the stochastic signal of cosmological origin is blended with strong astrophysical foregrounds, most notably the unresolved galactic binaries that form a confusion background and the contribution from extragalactic compact binary coalescences (CBCs) Tinto and Dhurandhar (2005); Smith et al. (2019); Barack and Cutler (2004); Caprini et al. (2024); Boileau et al. (2021a). Disentangling these different components—and thereby establishing a confident detection of the SGWB from a FOPT—requires dedicated statistical analysis.  [^1] A further challenge arises from deterministic, individually resolvable signals such as those from galactic binaries, massive black hole binaries, and extreme mass-ratio inspirals, which must be jointly analyzed along with stochastic components in the so-called global fit framework Cornish and Crowder (2005); Littenberg and Cornish (2023); Rosati and Littenberg (2024). This formidable task remains under active and will not be considered in this work.
@@
Significant progress has recently been achieved both in theoretical modeling and in data analysis techniques. The goal of this work is to bridge these two fronts—to integrate state-of-the-art theoretical developments with realistic experimental considerations—and to assess the achievable scientific outcomes in line with recent efforts within the community Gowling and Hindmarsh (2021); Gowling et al. (2023); Boileau et al. (2023); Caprini et al. (2024); Lewicki et al. (2024); Huang et al. (2025).
@@
For phenomenological studies, the interplay between GW observations and collider measurements has become a major focus of research Ramsey-Musolf (2020). In particular, the cubic and quartic self-couplings of the Higgs boson—central to understanding the scalar potential and the mechanism of electroweak symmetry breaking, but notoriously difficult to measure at colliders Di Vita et al. (2018); Zabinski (2023)—serve as key observables that highlight the complementarity between collider and GW approaches Alves et al. (2019).
@@
In this work, we demonstrate the above pipeline using the simplest extension of the Standard Model that includes a real scalar singlet—the xSM Barger et al. (2008); Profumo et al. (2007, 2015); Huang et al. (2017); Gonçalves et al. (2025). This study extends our previous phenomenological analyses Alves et al. (2019) by introducing a more realistic statistical treatment tailored to space-based GW experiments.
The remainder of this paper is organized as follows. In Sec. 2, we introduce the xSM and summarize its EWPT dynamics. Sec. 3 presents the GW spectrum from the EWPT, together with the astrophysical foreground and background models considered in this work. In Sec. 4, we describe the detector response, noise modeling, and data analysis procedures for space-based interferometers. Sec. 5 applies the statistical inference framework to the simulated data, maps the reconstructed spectral constraints onto the xSM parameter space, and derives the corresponding implications for Higgs self-couplings. Finally, Sec. 6 summarizes our results and discusses their implications for future GW missions and collider experiments. In Appendix A, we provide the detailed statistical inference framework, including a comparison between Fisher forecasts and Bayesian MCMC results.
@@ 2 The Model
In this section, we introduce the particle physics model and its parameter space, defined by five independent parameters. Once the model parameters are fixed, the phase transition parameters can be determined; these are the quantities that directly enter the prediction of the GW spectrum. Throughout this work, we follow the same convention as in our previous study Alves et al. (2019), to which we refer the reader for further details.
@@
The model “xSM” is the simplest extension of the Standard Model and has been extensively studied in the literature due to its simplicity and its ability to accommodate a strongly FOPT Barger et al. (2008); Profumo et al. (2007, 2015); Huang et al. (2017). This model is obtained by coupling the SM Higgs doublet with a real scalar gauge singlet, $S=v_{s}+s$, with the following potential
@@
with its vacuum expectation values $v_{\mathrm{EW}}$ and $G^{+},G^{0}$, the Goldstone modes. Through the two minimization conditions around the vacuum ($v_{\mathrm{EW}}$$v_{s}$), $\mu$ and $b_{2}$ are given by:
@@
These five parameters define the parameter space of the xSM model, which is constrained firstly by the theoretical requirement of the Higgs potential and by experimental measurements of various physical observables. The former includes the boundedness of the potential from below, electroweak vacuum stability at zero temperature, perturbativity, perturbative unitarity, etc. The latter includes precision measurements of the $W$ mass, the oblique electroweak corrections, the Higgs signal strengths, searches for a heavier Higgs-like scalar, etc. These have been extensively studied in the literature and our previous study Alves et al. (2019). Our study in this work is based on the phenomenologically viable parameter space, characterized by a large dataset of parameter space points that pass above the constraints. The readers are referred to the previous work for further details on the phenomenological constraints.
@@
$\begin{split}i\lambda_{h_{1}h_{1}h_{1}}&=6\big[\lambda vc_{\theta}^{3}+\tfrac{1}{4}c_{\theta}^{2}s_{\theta}(2a_{2}v_{s}+a_{1})+\tfrac{1}{2}a_{2}vc_{\theta}s_{\theta}^{2}\\ &+\tfrac{1}{3}s_{\theta}^{3}(3b_{4}v_{s}+b_{3})\big],\\ i\lambda_{h_{1}h_{1}h_{1}h_{1}}&=6(\lambda c_{\theta}^{4}+a_{2}s_{\theta}^{2}c_{\theta}^{2}+b_{4}s_{\theta}^{4}).\end{split}$ (2.7)
@@
We will later use $\delta\kappa_{3}$ and $\delta\kappa_{4}$ to assess how GW measurement can constrain Higgs self-couplings.
@@ 2.2 Phase Transition Parameters
For the GW spectrum generated by phase transitions, it is not the set of five model parameters in Eq. (2.6) that appears directly, but rather a derived set of phase transition parameters Caprini et al. (2024):
@@
which describes the thermodynamic properties of the EWPT. Specifically, $T_{\mathrm{n}}$ is the temperature at which the phase transition occurs, $\alpha$ quantifies the released vacuum energy normalized to the radiation energy density at $T_{\mathrm{n}}$, $\beta/H_{\mathrm{n}}$ characterizes the inverse time scale of the transition, and $v_{w}$ denotes the bubble wall velocity. These parameters are computed utilizing the most recent version of PhaseTracer with action fitting enabled. Bian et al. (2025); Athron et al. (2025, 2020).
In the Bayesian statistical analysis presented later, we do not directly infer the above thermodynamic parameters. Instead, our analysis is performed on a more direct spectral model characterized by two parameters: the overall amplitude $\Omega_{0}$ and the peak frequency $f_{\mathrm{p}}$. This choice avoids the issue of parameter degeneracy that would arise if one were to analyze the thermodynamic parameters directly, as will be discussed in subsequent sections. Once constraints on $\Omega_{0}$ and $f_{\mathrm{p}}$ are obtained, we then employ the mapping procedure introduced later to translate these results into constraints on the thermodynamic parameters and further into the particle-physics model parameters defined in Eq. (2.6). The detailed relations between these parameter sets, which require the calculation of the finite-temperature effective potential and the associated phase transition dynamics, can be found in our previous work Alves et al. (2019) and will not be repeated here.
Finally, we note that there remain significant theoretical uncertainties (see, e.g., Croon et al. (2021); Athron et al. (2023, 2024c); Lewicki et al. (2024); Zhu et al. (2025)) in determining the above phase transition parameters. These arise from several sources: the construction of the finite-temperature effective potential and the nucleation rate calculation, including the issue of gauge dependence for the perturbative calculation of the effective potential Patel and Ramsey-Musolf (2011); Arunasalam and Ramsey-Musolf (2022); Löfgren et al. (2023); Hirvonen et al. (2022); and the calculation of the parameters $T_{\mathrm{n}}$, $\alpha$, $\beta/H_{\mathrm{n}}$, and $v_{w}$, where different approaches with varying levels of rigor exist Guo et al. (2021b). For simplicity, we do not consider these theoretical uncertainties in the following statistical analysis, and we refer to Lewicki et al. (2024) for a study of this kind based on the FIM analysis.
@@ 3 Gravitational Waves: Cosmological and Astrophysical
While the target is the GW signal from the EWPT, one must also account for astrophysical and other possible cosmological backgrounds or foregrounds. These additional signals contribute incoherently to the observed data, thereby increasing the challenge of detecting the target signal. As a first step and for simplicity, we include the background from extragalactic CBCs, while a more complete treatment of other additions will follow the same methodology, involving only additional computational complexity.
@@ 3.1 Electroweak Phase Transition
In a FOPT, vacuum bubbles are nucleated, then grow, collide, and merge, driving the universe from a metastable to a stable state Guth and Weinberg (1981); Hindmarsh et al. (2021). The expansion and collision of these bubbles disturb the surrounding plasma in the form of sound waves (SWs) and magneto-hydrodynamic (MHD) turbulence Caprini and others (2016); Hindmarsh et al. (2014). As a result, the SGWB from a FOPT has three main contributions: (i) bubble collisions Kosowsky et al. (1992a, b); Kosowsky and Turner (1993); Huber and Konstandin (2008); Jinno and Takimoto (2017, 2019), (ii) SWs Hindmarsh et al. (2014, 2015); Cai et al. (2023); Roper Pol et al. (2024), and (iii) MHD turbulence Kosowsky et al. (2002); Caprini et al. (2009); Binetruy et al. (2012); Gogoberidze et al. (2007); Niksa et al. (2018); Roper Pol et al. (2020). However, the bubble-collision contribution is usually negligible compared to SWs for the EWPT, and that of the MHD turbulence is both subdominant and still uncertain. Thus, we consider only the dominant contribution from SWs.
The spectrum from SWs takes the following form Caprini and others (2016); Guo et al. (2021a):
@@
$v_{\mathrm{w}}\left(\frac{f}{f_{\mathrm{p}}}\right)^{3}\left(\frac{7}{4+3(f/f_{\mathrm{p}})^{2}}\right)^{7/2}\Upsilon\!\left(\tau_{\mathrm{sw}}\right)\;,$
where $H_{\mathrm{n}}$ is the Hubble rate at the nucleation temperature $T_{\mathrm{n}}$, $f_{\mathrm{p}}$ is the peak frequency Romero et al. (2021),
@@
and $\Upsilon$ is a suppression factor Guo et al. (2021a), a function of the finite lifetime Ellis et al. (2019) of the SWs $\tau_{\text{sw}}$, which in a radiation-dominated universe takes the following form:
@@
while the formula for a universe with a generic expansion rate was derived recently in Guo et al. (2024).
The lifetime $\tau_{\text{sw}}$ is usually chosen to be the time for the onset of MHD turbulence, $\tau_{\mathrm{sw}}=\frac{R_{\mathrm{pt}}}{\overline{U}_{f}}\;$ Weir (2018), where $R_{\mathrm{pt}}=(8\pi)^{1/3}v_{\mathrm{w}}/\beta$ is the mean bubble separation, and $\overline{U}_{f}=\sqrt{\frac{3\kappa_{\mathrm{sw}}\alpha}{4(1+\alpha)}}$ is the root-mean-square fluid velocity Weir (2018). For convenience, one can also define $K\equiv\frac{\alpha}{1+\alpha}$ Caprini et al. (2024), which encapsulates the dependence on transition strength.
@@
We set the spectral parameters to be log uniformly distributed in the ranges $\Omega_{0}\in(10^{-20},10^{-5})$ and $f_{\mathrm{p}}\in(10^{-5},1)\penalty 10000\ \mathrm{Hz}$, as the priors, taking into account the theoretical constraints of the xSM (as illustrated in Fig. 7) and following also Caprini et al. (2024); Romero et al. (2021). The corresponding ranges for phase transition parameters are: $\beta/H_{\mathrm{n}}\in(10^{-1},10^{5})$, $T_{\mathrm{n}}\in(1,10^{5})\penalty 10000\ \text{GeV}$, and $K\in(10^{-3},1)$, consistent with the sampling reconstruction results shown in Fig. 6. In addition, the wall velocity and the relativistic degrees of freedom are fixed to $v_{w}=1.0$ and $g_{*}=100$, respectively for simplicity, and the efficiency factor $\kappa_{\text{sw}}$ is taken as a function of $\alpha$ and $v_{w}$ (see the appendix of Espinosa et al. (2010) for the fitting formula). We note that choosing different values of the parameters such as $v_{w}$ leads to corresponding changes in the Bayesian inference, but not qualitatively on the conclusions obtained in this work.
@@
$v_{s}\simeq 34.73\ \mathrm{GeV},\quad m_{h_{2}}\simeq 573.59\ \mathrm{GeV},\quad\theta\simeq 0.099,\quad b_{3}\simeq-893.83\ \mathrm{GeV},\quad b_{4}\simeq 2.39,$
@@
$\frac{\beta}{H_{n}}=776.3,\quad T_{\mathrm{n}}=50.16\penalty 10000\ \mathrm{GeV},\quad K=0.289,$
@@
$\Omega_{0}=9.97\times 10^{-13},\quad f_{\mathrm{p}}=7.4\times 10^{-3}\penalty 10000\ \mathrm{Hz}.$